how to solve Three research departments have 7,9,8 and members, respectively. Each department is to select a delegate and an alternate to represent the department at a conference. In how many ways can this be done?
Answer by VFBundy(438)

7-member committee: 7*6 = 42 ways
8-member committee: 8*7 = 56 ways
9-member committee: 9*8 = 72 ways

Total ways: 42*56*72 = 169,344